# Superconductivity in Electron Liquids: Precision Many-Body Treatment of Coulomb Interaction

### v1

@@ I Introduction
Superconductivity (SC), a macroscopic quantum phenomenon with far-reaching fundamental physics and technological implications, is a focal point of condensed matter research. For the first half century after its discovery, lack of a theoretical basis for understanding the physics of interacting electrons inhibited the development of a theory of superconductivity. Following the work of Fröhlich [frohlich1950] and the experimental confirmation by the Rutgers and NBS groups [reynolds1950, maxwell1950], the importance of the electron phonon interaction was recognized, but how it competed with Coulomb repulsion was not clear. Bardeen, Cooper and Schrieffer (BCS) then provided our basic understanding of superconductivity in terms of a phonon-mediated attraction leading to a pairing instability [cooper_bound_1956-1, bardeen_theory_1957, bardeen_microscopic_1957]. The original BCS paper took the radical step of ignoring Coulomb effects altogether but among many other important concepts introduced the idea that if a weak pairing interaction existed at some energy scale, its importance increased as the scale was lowered. This meant that superconductivity should be understood as a low energy instability of the electron gas and that the competition was between a phonon-mediated interaction and an effective electron-electron repulsion defined at the scale of the phonon frequency. This insight motivated studies of the effective Coulomb interaction in the pairing (Cooper) channel at low energy, obtained by “downfolding” the fundamental Coulomb interaction. The first downfolding argument was provided by Bogoliubov, Tolmachev, and Shirkov [bogoliubov59, tolmachev1961logarithmic] who showed that the same Cooper logarithm that leads to SC for attractive interactions will renormalize a frequency independent local Coulomb repulsion to the weak coupling limit at low energies. The resulting dimensionless parameter, $\mu^{*}>0$ describing the effect of Coulomb interactions in the pairing channel at the scale of the phonon frequency, was later named the “pseudopotential” by Morel and Anderson [morel_anderson]).
@@
The concepts of phonon-mediated attraction and renormalized Coulomb repulsion are encoded in the Migdal-Eliashberg (ME) theory [migdal1958, eliashberg1960interactions, eliashberg1961temperature], which goes beyond the original BCS theory by considering both the dynamic electron-phonon (e-ph) and electron-electron (e-e) interactions and provides the foundational framework for understanding phonon-mediated superconductors. ME theory as conventionally implemented relies on a “downfolding” approximation [eliashberg1960interactions, eliashberg1961temperature, tolmachev1961logarithmic, mcmillan1968, allen1975, EPW1, EPW2, AEPW] that takes advantage of the large energy scale separation between a typical Fermi energy, $E_{F}$, and a typical phonon frequency $\omega_{D}$ to project the equations into the low-energy subspace defined by an effective Fermi-surface theory of quasiparticles, replacing bare e-ph interactions with screened effective e-ph coupling, $\lambda$ [eliashberg1960interactions, eliashberg1961temperature], and assuming that Coulomb interactions effects reduce to Landau parameters, along with the pairing channel instantaneous pseudopotential $\mu^{*}$. It is important to note that although the adiabatic limit (when $\omega_{D}/E_{F}\ll 1$) justifies the formulation in terms of a downfolded theory, as it is currently implemented many aspects of the ME theory remain semi-phenomenological. Coulomb-induced quasiparticle renormalizations are neglected, as are possible renormalizations of the electron-phonon coupling by Coulomb fluctuations as well as non-local effects coming from a scale dependence of screening, particularly pronounced in two dimensional materials [Simonato23]. The Coulomb pseudopotential $\mu^{*}$ is chosen in an ad-hoc or phenomenological manner  [RAINER1986371]. This creates critical challenges both for first-principles predictions of transition temperatures and for quantifying unconventional, i.e. not involving e-ph coupling, pairing mechanisms.
@@
Perhaps the most controversial issue is the magnitude and sign of $\mu^{*}$. The first theoretical estimates were based on the unphysical approximation that Coulomb interactions are screened at all frequencies. When this approximation was subsequently removed by employing the "random phase" approximation (RPA) to compute the more physical dynamically screened interaction, Takada [takada1978plasmon, takada1993] and Rietschel and Sham [rietschel] found that the dynamical effects of the Coulomb potential may lead to attractive $\mu^{*}<0$ and s-wave superconductivity even in the absence of phonons, if the electron-electron interaction is strong enough. Recent studies again within the RPA, now extended to all paring channels [23tao, xiansheng, pengcheng] confirm the validity of this theoretical conclusion.
@@
The RPA-based finding of purely electronically mediated superconductivity is inconsistent with a large body of empirical evidence, and is also internally inconsistent. For a recent discussion see Ref. [dassarma2025]. In the electron gas, the strength of the Coulomb interaction is parametrized by the dimensionless Wigner-Seitz radius $r_{s}=\left(\frac{9\pi}{4}\right)^{1/3}\frac{me^{2}}{\hbar^{2}k_{\text{F}}}$, in effect the ratio of interaction to kinetic energies at the scale of the interelectron spacing. The RPA is a good approximation for $r_{s}\lesssim 1$ while the Coulombic pairing was found at $r_{s}\gtrsim 2$ where beyond-RPA effects, including vertex corrections and renormalization of the single-particle propagators [RPA_mu2, 23tao], may be important.
@@
Due to these theoretical uncertainties, $\mu^{*}$ is often treated as an adjustable parameter, with values empirically chosen between 0.1 and 0.2 to reconcile theory with experiments. This range leads to large (sometime orders of magnitude) uncertainty when predicting $T_{c}$ in materials where SC depends on a delicate balance between phonon-mediated attraction and $\mu^{*}$. For instance, ME theory predictions for $T_{c}$ in aluminum, a workhorse of superconducting electronics for transition-edge sensors [irwin_transition-edge_2005] and qubits [qubit_poisoning], deviate from experiments by a factor of two [pellegrini2024ab], while for elemental lithium the theory overestimates $T_{c}$ by three orders of magnitude [ashcroft1, hubbard_mu, tuoriniemi_superconductivity_2007]. These discrepancies severely limit the possibility of reliable predictive design for quantum devices requiring precise control of energy gaps and thermal noise. Even more strikingly, some transition-metal based compounds (e.g., V, Nb3Sn [RevModPhys.62.1027]), alkali-doped picene [PhysRevB.84.020508], and high-pressure hydrides [kostrzewa_anomalously_2018] are best described by $\mu^{*}$ values from $0.2$ to $0.5$, well outside of the limits following from the static [morel_anderson] and dynamic RPA [margine2016electron, RPA_mu, RPA_mu2]. While more advanced theories such as the T-matrix approximation [scalapino1969electron, gladstone2018superconductivity] have been proposed to address the discrepancy, their development rather underscores the need for a systematic precise microscopic theory of $\mu^{*}$ as opposed to the case-selective use of uncontrolled approximations.
@@
Equally important is accurate treatment of the effective coupling, $\lambda$, which quantifies the phonon-mediated attraction between quasiparticles at the Fermi surface within the downfolding procedure. State-of-the-art ab initio methods, such as the density functional perturbation theory (DFPT) [giustino_2017, EPW1, EPW2, AEPW, pellegrini2024ab], calculate $\lambda$ from the ground-state energy response to lattice distortions, an approximation validated for weakly correlated superconductors [RAINER1986371]. However, it remains unknown how accurate this proxy is for correlated systems, where strong renormalization effects are expected to alter effective e-ph interactions and thus SC [RPA_mu, RPA_mu2, zfactor_2022, 23tao]. Removing this uncertainty is crucial for extending the predictive power of such a method as density functional theory plus dynamical mean-field theory (DFT+DMFT) [dftdmft1, dftdmft2, dftdmft3, dftdmft4, dftdmft5, dftdmftreview, dmft_sc1, dmft_sc2] to strongly correlated superconductors.
@@
The downfolded ME framework also faces conceptual challenges on top of practical difficulties with precise evaluation of $\mu^{*}$ and $\lambda$. The conventional approach becomes inadequate in the strong e-ph coupling regime, e.g. in high-pressure hydrides with large quantum nuclear effects [kostrzewa_anomalously_2018], near structural transitions, or due to the formation of bipolarons [bipolaron1, bipolaron2, bipolaron3].
@@
These limitations have motivated efforts to transcend the downfolding procedure altogether. Richardson and Ashcroft suggested solving the system of ME equations “as is” with full momentum-frequency dependence of both the e-ph and screened (at the RPA level) e-e interactions [ashcroft1, ashcroft2]. Their approach successfully predicted $T_{c}$ for lithium [tuoriniemi_superconductivity_2007], but its reliance on the RPA casts doubt on its applicability for electron densities relevant to most metals. This and the large computational cost have prevented this method from becoming widely adopted (but see [veld23] for recent applications in two dimensions). Superconducting DFT [SCDFT1, SCDFT2, profeta_superconductivity_2006] offers an alternative by generalizing DFT to include the superconducting order parameter, with screening effects approximated by the RPA-based ansatz. Recently, it was adapted within the ME framework [scdft-eliashberg], offering a potentially fruitful avenue for future research. However, such issues as contributions coming from spin fluctuations [kawamura_benchmark_2020] and vertex corrections to the effective coupling remain unresolved, and the method as implemented to date relies on the RPA. To our knowledge, a consistent and quantitatively accurate microscopic treatment of Coulomb effects in the theory of SC has not yet been achieved.
@@
Importantly, we also provide limits of validity of the theory, which fails at extremely high electron densities when the plasmon frequency, $\omega_{\text{p}}$, softens, in 2D when the plasmon mode is no longer gapped, in materials with soft collective excitations emerging from strong correlations, and systems with low conduction electron density, revealing regimes where the dynamic nature of screening is most relevant. While 3D electron densities corresponding to $\omega_{\text{p}}\ll E_{\text{F}}$ are beyond current experimental reach terrestrially, they can be found in ultra-dense astrophysical objects [trubnikov1968white, ginzburg_superconductivity_1968].
As a critical application, we compute Coulomb pseudopotentials for the uniform electron gas (UEG) using recently developed variational diagrammatic Monte Carlo (DiagMC) method [diagmc1, diagmc2, diagmc2010, detdiagmc, diagmclatest, diagmckozik, vdiagmc1, vdiagmc2]. These results provide precise parameterization of $\mu^{*}$ as a function of electron density to be used in ME treatment of real materials. Our estimates of the “bare” pseudopotential, obtained by first computing the quasiparticle scattering amplitude at the Fermi surface and then renormalizing it via the Cooper/Tolmachev logarithm, are significantly larger than estimates based on static screening or RPA in the moderate-density regime, bridging a crucial gap between theory and experiment. We also compute the quasiparticle e-ph vertex for the UEG, finding that DFPT yields remarkably accurate electron-phonon coupling values for simple metals.
By combining these advances with calculations of the precursory Cooper flow (PCF) of the anomalous vertex function in the normal state [pengcheng], we propose an ab initio workflow, capable of predicting low-temperature SC and quantum phase transition points well beyond the reach of conventional methods. With this approach, we revisited the problem of superconducting $T_{\text{c}}$ in various simple metals under ambient and high-pressure conditions. Our method yields orders-of-magnitude improvements in predicting $T_{\text{c}}$ for sub-Kelvin superconductors compared to previous estimates based on a phenomenological $\mu^{*}$. We find that Mg and Na are close to a normal-SC quantum phase transition, offering an opportunity to study quantum critical scaling below 1010K. We also predict that Al will undergo a transition to the normal state under pressure exceeding 6060GPa. In the UEG context, this work establishes a complete and controlled first-principles framework for understanding the interplay between the electron correlations and e-ph interactions in SC. If rigorously followed in simulations of real materials, it will radically improve our ability to design next-generation superconducting materials.
@@ II.1 Electron-Phonon Problem
Our goal is to address electron-phonon SC in crystals using effective field theory (EFT) [polchinski_effective_1999] derived on the assumption that the electron mass $m$ is much less than an ionic mass $M$. The mass difference has three essential consequences: the typical phonon frequency $\omega_{\text{D}}$ is suppressed relative to $E_{\text{F}}$ by a factor of $(m/M)^{1/2}$, ensuring that electrons adiabatically adjust to the ionic motion; the momentum transferred by an electron to an ion in collision is very small, justifying a linearization of the electron-ion coupling; and the separation of spatial and temporal scales of electronic and phononic physics justifies a controlled EFT treatment–in particular, once properties of the pure Coulomb system (no phonons) are established at energies greater than $\omega_{\text{D}}$, further corrections to the e-ph vertex are suppressed by $(m/M)^{1/2}$ allowing for simple perturbative analysis of the low-energy theory.
@@
where $\psi$, $\bar{\psi}$ are Grassmann fields for electrons, and $u$ is the ionic displacement field rescaled by $\sqrt{M}$ (this standard procedure applies to any number of ions in the unit cell). Here, $S_{\text{e}}$ stands for full many-electron action without any approximations. We distinguish our theory from the conventional DFT approach, which replaces $S_{\text{e}}$ with a non-interacting reference Hamiltonian governed by an approximate exchange-correlation potential [kohn1964, kohn1965]. The DFT approach does not capture the dynamical effects of electron-electron interactions.
@@ II.2 Bethe-Salpeter Equation (BSE)
It consists of two terms with the first one, ${\Sigma}_{\mathbf{k}\omega}^{\text{e}}$, coming entirely from the non-perturbative Coulomb interactions. The second term has the structure of the self-consistent Fock diagram based on the dressed phonon-mediated interaction $W^{\text{ph}}$ shown in Fig. 2. By construction, $W^{\text{ph}}$ incorporates all non-perturbative Coulomb effects such as screening and vertex corrections. Migdal’s theorem ensures that vertex corrections higher-order in $W^{\text{ph}}$ are suppressed by $\omega_{\text{D}}/E_{\text{F}}$, justifying truncation at this level of accuracy [migdal1958]. Similarly, the anomalous (superconducting) self-energy has a Coulomb contribution, ${\Sigma}_{\mathbf{k}\omega}^{\text{a},\text{e}}$, and a self-consistent “Fock-type” contribution based on the phonon-mediated interaction
@@
These two equations are in effect the standard Migdal-Eliashberg equations, expressed in the current notations. The transition temperature is typically found by linearizing the second equation. Analysis even of the linearized equation can be challenging for low $T_{c}$ superconductors where a dense computational mesh is required and for strong coupling situations, where extrapolation of the leading eigenvalues of the linearized gap equation is not straightforward. If the goal is limited to prediction of $T_{\text{c}}$, then following Ref. [pengcheng], it is more efficient to probe SC from the high-temperature normal state ($T\gg T_{\text{c}}$) by solving Bethe-Salpeter equation (BSE) for the anomalous vertex function defined from the linear response of the anomalous self energy to an applied pairing field. Here, we apply this approach in the context of the electron-phonon SC by considering response to an infinitesimal symmetry-breaking pair-field term ($\eta^{\text{a}}\to 0$):

### v2

@@ I Introduction
Superconductivity (SC), a macroscopic quantum phenomenon with far-reaching fundamental physics and technological implications, is a focal point of condensed matter research. For the first half century after its discovery, lack of a theoretical basis for understanding the physics of interacting electrons inhibited the development of a theory of superconductivity. Following the work of Fröhlich [26] and the experimental confirmation by the Rutgers and NBS groups [84, 65], the importance of the electron phonon interaction was recognized, but how it competed with Coulomb repulsion was not clear. Bardeen, Cooper and Schrieffer (BCS) then provided our basic understanding of superconductivity in terms of a phonon-mediated attraction leading to a pairing instability [21, 8, 7]. The original BCS paper took the radical step of ignoring Coulomb effects altogether but among many other important concepts introduced the idea that if a weak pairing interaction existed at some energy scale, its importance increased as the scale was lowered. This meant that superconductivity should be understood as a low energy instability of the electron gas and that the competition was between a phonon-mediated interaction and an effective electron-electron repulsion defined at the scale of the phonon frequency. This insight motivated studies of the effective Coulomb interaction in the pairing (Cooper) channel at low energy, obtained by “downfolding” the fundamental Coulomb interaction. The first downfolding argument was provided by Bogoliubov, Tolmachev, and Shirkov [13, 103] who showed that the same Cooper logarithm that leads to SC for attractive interactions will renormalize a frequency independent local Coulomb repulsion to the weak coupling limit at low energies. The resulting dimensionless parameter, $\mu^{*}>0$ describing the effect of Coulomb interactions in the pairing channel at the scale of the phonon frequency, was later named the “pseudopotential” by Morel and Anderson [69]).
@@
The microscopic description of phonon-mediated attraction is encoded in the Migdal-Eliashberg theory, which goes beyond the original BCS theory by rigorously treating the dynamic electron-phonon (e-ph) interaction and provides the foundational framework for understanding phonon-mediated superconductors [68, 22, 23]. While the original formulation [68, 22] focused on the e-ph interaction, subsequent developments [103, 69] incorporated Coulomb repulsion via the static pseudopotential $\mu^{*}$ and following modern convention we refer to this combined theory as “Migdal-Eliashberg” or ME theory. ME theory as conventionally implemented relies on a “downfolding” approximation [22, 23, 103, 67, 4, 32, 78, 61] that takes advantage of the large energy scale separation between a typical Fermi energy, $E_{\text{F}}$, and a typical phonon frequency $\omega_{\text{D}}$ to project the equations into the low-energy subspace defined by an effective Fermi-surface theory of quasiparticles, replacing bare e-ph interactions with screened effective e-ph coupling, $\lambda$ [22, 23]. It is important to note that although the adiabatic limit (when $\omega_{\text{D}}/E_{\text{F}}\ll 1$) justifies the formulation in terms of a downfolded theory, as it is currently implemented many aspects of the ME theory remain semi-phenomenological. Coulomb-induced quasiparticle renormalizations are neglected, as are possible renormalizations of the electron-phonon coupling by Coulomb fluctuations as well as non-local effects coming from a scale dependence of screening, particularly pronounced in two dimensional materials [96]. The Coulomb pseudopotential $\mu^{*}$ is chosen in an ad-hoc or phenomenological manner  [83]. This creates critical challenges both for first-principles predictions of transition temperatures and for quantifying unconventional, i.e. not involving e-ph coupling, pairing mechanisms.
@@
Perhaps the most controversial issue is the magnitude and sign of $\mu^{*}$. The first theoretical estimates were based on the unphysical approximation that Coulomb interactions are screened at all frequencies. When this approximation was subsequently removed by employing the “random phase” approximation (RPA) to compute the more physical dynamically screened interaction, Takada [100, 101] and Rietschel and Sham [89] found that the dynamical effects of the Coulomb potential may lead to attractive $\mu^{*}<0$ and s-wave superconductivity even in the absence of phonons, if the electron-electron interaction is strong enough. Recent studies again within the RPA, now extended to all paring channels [108, 15, 44] confirm the validity of this theoretical conclusion.
@@
The RPA-based finding of purely electronically mediated superconductivity is inconsistent with a large body of empirical evidence, and is also internally inconsistent. For a recent discussion see Ref. [92]. In the electron gas, the strength of the Coulomb interaction is parametrized by the dimensionless Wigner-Seitz radius $r_{\text{s}}=\left(\frac{9\pi}{4}\right)^{1/3}\frac{me^{2}}{\hbar^{2}k_{\text{F}}}$, in effect the ratio of interaction to kinetic energies at the scale of the interelectron spacing. The RPA is a good approximation for $r_{\text{s}}\lesssim 1$ while the Coulombic pairing was found at $r_{\text{s}}\gtrsim 2$ where beyond-RPA effects, including vertex corrections and renormalization of the single-particle propagators [2, 108], may be important.
@@
Due to these theoretical uncertainties, $\mu^{*}$ is often treated as an adjustable parameter, with values empirically chosen between 0.1 and 0.2 to reconcile theory with experiments. This range leads to large (sometime orders of magnitude) uncertainty when predicting $T_{c}$ in materials where SC depends on a delicate balance between phonon-mediated attraction and $\mu^{*}$. For instance, ME theory predictions for $T_{c}$ in aluminum, a workhorse of superconducting electronics for transition-edge sensors [47] and qubits [6], deviate from experiments by a factor of two [72], while for elemental lithium the theory overestimates $T_{c}$ by three orders of magnitude [86, 11, 105]. These discrepancies severely limit the possibility of reliable predictive design for quantum devices requiring precise control of energy gaps and thermal noise. Even more strikingly, some transition-metal based compounds (e.g., V, Nb3Sn [16]), alkali-doped picene [99], and high-pressure hydrides [54] are best described by $\mu^{*}$ values from $0.2$ to $0.5$, well outside of the limits following from the static [69] and dynamic RPA [62, 91, 2]. While more advanced theories such as the T-matrix approximation [93, 34] have been proposed to address the discrepancy, their development rather underscores the need for a systematic precise microscopic theory of $\mu^{*}$ as opposed to the case-selective use of uncontrolled approximations.
@@
Equally important is accurate treatment of the effective coupling, $\lambda$, which quantifies the phonon-mediated attraction between quasiparticles at the Fermi surface within the downfolding procedure. State-of-the-art ab initio methods, such as the density functional perturbation theory (DFPT) [33, 32, 78, 61, 72], calculate $\lambda$ from the ground-state energy response to lattice distortions, an approximation validated for weakly correlated superconductors [83]. However, it remains unknown how accurate this proxy is for correlated systems, where strong renormalization effects are expected to alter effective e-ph interactions and thus SC [91, 2, 3, 108]. Removing this uncertainty is crucial for extending the predictive power of such a method as density functional theory plus dynamical mean-field theory (DFT+DMFT) [5, 55, 40, 10, 70, 71, 52, 48] to strongly correlated superconductors.
@@
The downfolded ME framework also faces conceptual challenges on top of practical difficulties with precise evaluation of $\mu^{*}$ and $\lambda$. The conventional approach becomes inadequate in the strong e-ph coupling regime, e.g. in high-pressure hydrides with large quantum nuclear effects [54], near structural transitions, or due to the formation of bipolarons [24, 20, 109].
@@
These limitations have motivated efforts to transcend the downfolding procedure altogether. Richardson and Ashcroft suggested solving the system of ME equations “as is” with full momentum-frequency dependence of both the e-ph and screened (at the RPA level) e-e interactions [86, 88]. Their approach successfully predicted $T_{c}$ for lithium [105], but its reliance on the RPA casts doubt on its applicability for electron densities relevant to most metals. This and the large computational cost have prevented this method from becoming widely adopted (but see [46] for recent applications in two dimensions). Superconducting DFT [60, 63, 79] offers an alternative by generalizing DFT to include the superconducting order parameter, with screening effects approximated by the RPA-based ansatz. Recently, it was adapted within the ME framework [1], offering a potentially fruitful avenue for future research. However, such issues as contributions coming from spin fluctuations [50] and vertex corrections to the effective coupling remain unresolved, and the method as implemented to date relies on the RPA. To our knowledge, a consistent and quantitatively accurate microscopic treatment of Coulomb effects in the theory of SC has not yet been achieved.
@@
Importantly, we also provide limits of validity of the theory, which fails at extremely high electron densities when the plasmon frequency, $\omega_{\text{p}}$, softens, in 2D when the plasmon mode is no longer gapped, in materials with soft collective excitations emerging from strong correlations, and systems with low conduction electron density, revealing regimes where the dynamic nature of screening is most relevant. While 3D electron densities corresponding to $\omega_{\text{p}}\ll E_{\text{F}}$ are beyond current experimental reach terrestrially, they can be found in ultra-dense astrophysical objects [104, 31].
As a critical application, we compute Coulomb pseudopotentials for the uniform electron gas (UEG) using recently developed variational diagrammatic Monte Carlo (DiagMC) method [80, 81, 107, 90, 106, 56, 18, 39]. These results provide precise parameterization of $\mu^{*}$ as a function of electron density to be used in ME treatment of real materials. Our estimates of the “bare” pseudopotential, obtained by first computing the quasiparticle scattering amplitude at the Fermi surface and then renormalizing it via the Cooper/Tolmachev logarithm, are significantly larger than estimates based on static screening or RPA in the moderate-density regime, bridging a crucial gap between theory and experiment. We also compute the quasiparticle e-ph vertex for the UEG, finding that DFPT yields remarkably accurate electron-phonon coupling values for simple metals.
By combining these advances with calculations of the precursory Cooper flow (PCF) of the anomalous vertex function in the normal state [44], we propose an ab initio workflow, capable of predicting low-temperature SC and quantum phase transition points well beyond the reach of conventional methods. With this approach, we revisited the problem of superconducting $T_{\text{c}}$ in various simple metals under ambient and high-pressure conditions. Our method yields orders-of-magnitude improvements in predicting $T_{\text{c}}$ for sub-Kelvin superconductors compared to previous estimates based on a phenomenological $\mu^{*}$. We find that Mg and Na are close to a normal-SC quantum phase transition, offering an opportunity to study quantum critical scaling below 1010K. We also predict that Al will undergo a transition to the normal state under pressure exceeding 6060GPa. In the UEG context, this work establishes a complete and controlled first-principles framework for understanding the interplay between the electron correlations and e-ph interactions in SC. If rigorously followed in simulations of real materials, it will radically improve our ability to design next-generation superconducting materials.
@@ II.1 Electron-Phonon Problem
Our goal is to address electron-phonon SC in crystals using effective field theory (EFT) [77] derived on the assumption that the electron mass $m$ is much less than an ionic mass $M$. The mass difference has three essential consequences: the typical phonon frequency $\omega_{\text{D}}$ is suppressed relative to $E_{\text{F}}$ by a factor of $(m/M)^{1/2}$, ensuring that electrons adiabatically adjust to the ionic motion; the momentum transferred by an electron to an ion in collision is very small, justifying a linearization of the electron-ion coupling; and the separation of spatial and temporal scales of electronic and phononic physics justifies a controlled EFT treatment–in particular, once properties of the pure Coulomb system (no phonons) are established at energies greater than $\omega_{\text{D}}$, further corrections to the e-ph vertex are suppressed by $(m/M)^{1/2}$ allowing for simple perturbative analysis of the low-energy theory.
@@
where $\psi$, $\bar{\psi}$ are Grassmann fields for electrons, and $u$ is the ionic displacement field rescaled by $\sqrt{M}$ (this standard procedure applies to any number of ions in the unit cell). Here, $S_{\text{e}}$ stands for full many-electron action without any approximations. We distinguish our theory from the conventional DFT approach, which replaces $S_{\text{e}}$ with a non-interacting reference Hamiltonian governed by an approximate exchange-correlation potential [41, 53]. The DFT approach does not capture the dynamical effects of electron-electron interactions.
@@ II.2 Bethe-Salpeter Equation (BSE)
It consists of two terms with the first one, ${\Sigma}_{\mathbf{k}\omega}^{\text{e}}$, coming entirely from the non-perturbative Coulomb interactions. The second term has the structure of the self-consistent Fock diagram based on the dressed phonon-mediated interaction $W^{\text{ph}}$ shown in Fig. 2. By construction, $W^{\text{ph}}$ incorporates all non-perturbative Coulomb effects such as screening and vertex corrections. Migdal’s theorem ensures that vertex corrections higher-order in $W^{\text{ph}}$ are suppressed by $\omega_{\text{D}}/E_{\text{F}}$, justifying truncation at this level of accuracy [68]. Similarly, the anomalous (superconducting) self-energy has a Coulomb contribution, ${\Sigma}_{\mathbf{k}\omega}^{\text{a},\text{e}}$, and a self-consistent “Fock-type” contribution based on the phonon-mediated interaction
@@
These two equations are in effect the standard Migdal-Eliashberg equations, expressed in the current notations. The transition temperature is typically found by linearizing the second equation. Analysis even of the linearized equation can be challenging for low $T_{c}$ superconductors where a dense computational mesh is required and for strong coupling situations, where extrapolation of the leading eigenvalues of the linearized gap equation is not straightforward. If the goal is limited to prediction of $T_{\text{c}}$, then following Ref. [44], it is more efficient to probe SC from the high-temperature normal state ($T\gg T_{\text{c}}$) by solving Bethe-Salpeter equation (BSE) for the anomalous vertex function defined from the linear response of the anomalous self energy to an applied pairing field. Here, we apply this approach in the context of the electron-phonon SC by considering response to an infinitesimal symmetry-breaking pair-field term ($\eta^{\text{a}}\to 0$):